# Foveated Retinotopy Improves Classification and Localization in Convolutional Neural Networks

### v4

@@ Results
We retrain pre-trained ResNet (He et al., 2015) networks on different variants of the ImageNet-1K dataset (Russakovsky et al., 2015) (see Methods “Datasets and Dataset Transformations”), either using a simple circular mask applied on the raw images (hereafter called the Cartesian dataset) or using log-polar transformed images (hereafter called the “retinotopic” dataset). Two variants of the training sets are also considered. In a first case (called the “regular” case), the mask or the log-polar transformed images are applied to the “regular” images. In a second case (called the “focus” case), the mask or the log-polar transformation is focused on the center of the bounding boxes surrounding the objects of interest and which are provided with the data set. The pre-trained networks are first re-trained on the “regular” version of the dataset, generating a first generation of networks, and then a second generation is trained by fine-tuning these networks on the “focus” dataset (see Figure 11). We then compare the classification accuracy of the original and the different re-trained networks. In this section we choose to express the accuracy as the percentage of correct categorization on the corresponding validation dataset.
@@
This difference between the two types of networks is even more manifest in Figure 2-B, where we compute the average accuracy over the test dataset for each single rotation angle. We first observe that the Cartesian networks show a decrease in accuracy with respect to the angle of rotation (with a symmetry with respect to horizontal flips), with a significant drop around $160^{\circ}$, “raw” and “focus” accuracies were degraded to $66.8\%$ and $68.1\%$, i.e. a drop of $16.1\%$ and $13.4\%$ ($64.3\%$ in “regular”, i.e. a drop of $15.8\%$). Strikingly, this effect is nearly absent in our retinotopic networks (see Figure 2-B), which show a flat (invariant) accuracy over the whole range of rotation angles, with a minimal degraded accuracy at $76.4\%$, i.e. a drop of $0.8\%$ ($71.6\%$ in “regular”, i.e. a drop of $2.9\%$). This marked difference can be interpreted as a consequence of the horizontal translation invariance found in classical CNNs. When applied to the retinotopic input space, this invariance transforms seamlessly into rotation invariance in visual space (Araujo and Dias, 1997) (see (Hao et al., 2021) for a proof). Note that for all networks, no maximum accuracy exceeds the value obtained without rotation (i.e. $0^{\circ}$ rotation angle) by more than $0.01\%$.
@@
To quantify the contribution of this sensitivity to translation, we consider here a new task, i.e. the visual search task, in which a visual object (of which the label is known in advance) needs to be localized over the entire image. Is is for instance known that such task allows for the quick retrieval of an image label (Potter, 1975). We design a protocol for each network to allow us to compare different visual shifts, each one corresponding to a potential fixation point, and to generate a map of the expected (or actual) accuracy as a function of the fixation point (see Methods “Visual object localization : Protocol”). In this protocol, a set of $11\times 11$ fixation points is defined (with the coordinate $(5,5)$ being the center of the image), and at each coordinate of the grid, a likelihood value is computed for the label of interest (see figure 13). In practice, a likelihood is given at each location from a softmax calculation over the different labels, as in the classical ResNet classifier, providing a value between 0 and $1$ for the label of interest. This projects the network output onto a $11\times 11$ Bernoulli probability space corresponding to the likelihood of detecting the given label at each position, finally providing a “heat map” on our $11\times 11$ grid in a way that is compatible with other localization protocols. We tested our likelihood protocol on the “regular” validation dataset (see Fig 5 for some examples).
@@
ImageNet-1K provides rich semantic links that allow the construction of task-specific datasets. It has previously been demonstrated that the use of fine-tuning to re-train networks such as Vgg16 (Simonyan and Zisserman, 2015) allows them to be applied to different tasks using the semantic network underlying ImageNet’s labels. Furthermore, it has been shown that the probability of a trained network performing a novel task (such as categorizing an animal) can be predicted using this semantic network, which links the outputs of a pre-trained network to a library of labels. This approach has been shown to be an effective method for learning to predict the presence of animals in images (Jérémie and Perrinet, 2023). We can then exploit the semantic connection underlying the ImageNet labels by using these networks to perform a categorization task on the second dataset, the Animal 10k dataset, as referenced by Yu et al. (2021), with the goal of identifying the animals depicted in the images. This dataset contains $10,015$ images, each depicting an animal. All images were integrated in a validation dataset to test the networks. The output of each network is interpreted as “animal” if the label represented by the highest likelihood after softmax is indeed an animal. We use this prediction to calculate the accuracy of predicting an animal for all networks. Note that the accuracy is slightly biased due to the absence of distractors, it only represents the probability of finding an animal when there is actually an animal in the input. Figure 8 (right panel) compares the accuracy improvement for both networks. The general accuracy is high (around $95\%$ in all cases) due to our simpler binary (animal/non-animal) classification task. Saccade improvement is still significant, with all networks reaching a maximal accuracy of $100\%$ after a saccade to the maximum likelihood position for the label “animal”. The highest improvement is once again found in the retinotopic case (“raw” $4.5\%$, Cartesian $4.1\%$, retinotopic $6.3\%$).
@@ Discussion
The first and main result of this study is to demonstrate the excellent ability of standard deep CNNs to deal with foveated retinotopic inputs, even though this mapping enforces a radical transformation of the visual inputs. ResNet networks easily adapt to these inputs and the accuracy rates achieved with retinotopic inputs are equivalent to those of the original models. This is surprising given that the networks used in this re-training process were previously trained on Cartesian images. Images with a log-polar transformation show a high degree of distortion, particularly a high compression of visual information around the fixation point and a degradation of textures in the periphery (see Figure 1). One possible hypothesis is that the degradation of texture during the frame of reference change may cause the network to rely on shape rather than texture (da Costa et al., 2024). A further study could entail a comparison of these visual processing parameters with those of human vision, thereby providing insight into their evolutionary trajectory.
@@
In addition, the log-polar transformation has the advantage of better invariance to zooms and rotations. First, this study shows that for the original ResNet network on “regular” images, the average accuracy with a zoom or a rotation dropped sharply compared to baseline accuracy, confirming that these simple geometrical transformations misled the networks. For rotations, the decline was steepest around $160^{\circ}$, demonstrating limited rotational invariance compared to humans (Rousselet et al., 2003; Guyonneau et al., 2006). The robustness differs slightly when studying the impact of zooms. This may be attributed to the transformation under examination, which already provides a zooming-in effect of the fovea that accentuates the information in the area surrounding the point of fixation. To test this hypothesis, we could train a retinotopic version without applying the logarithmic function to the eccentricity axis. Finally, this invariance to rotations and zooms comes at the cost of reduced invariance to translations. For images not centered on the region of interest, one would need to shift the fixation point to the region of interest, similar to eye saccades in biological vision. The effects of this trade-off can be observed in the comparison between the “regular” version of the networks and the “focus” version.
@@
In addition, the log–polar transformation confers improved invariance to zoom and rotation. In the original ResNet trained on “regular” Cartesian images, accuracy declined sharply under zooms or rotations, confirming that even simple geometric transformations can mislead the network. The steepest drop occurred around $160^{\circ}$, highlighting limited rotational invariance compared to humans (Rousselet et al., 2003; Guyonneau et al., 2006). Robustness to zoom was less affected, likely because the transformation itself magnifies foveal input relative to the periphery. This robustness, however, comes at the cost of reduced tolerance to translations: images misaligned with the fixation point require shifting the sampling center, analogous to saccades in biological vision. This trade-off is evident when comparing the “regular” and “focus” network variants. Similar effects have been reported previously:  Pun and Lee showed that log–polar wavelet signatures enhance joint scale–rotation invariance, while  Amorim et al. demonstrated improved rotational robustness in natural image categorization and highlighted the positional sensitivity introduced by inhomogeneous log–polar sampling. Finally, previous research has demonstrated that improved invariance to scale variations is often associated with greater robustness to input compression (Remmelzwaal et al., 2019). This dual robustness — the ability to maintain performance across scaled inputs and compressed representations — facilitates the application of categorisation models to high-resolution images without compromising accuracy. These properties are particularly valuable in computational neuroscience, where visual processing systems must generalise across diverse input conditions.
@@
One hypothesis is that the foveated retinotopic mapping implemented in the log-polar transform provides an efficient prior for visual object geometry. Indeed, a main source of variability in the view of an object comes from displacements of the observer relative to it, for example as the object scales in the visual field as one approaches it, or as the object rotates with a rotation of the head. The log-polar mapping allows for a more invariant representation which explicitly implements this prior for these displacements. Additionally, imposing translation invariance of the representation establishes a prior on the possible representations allowed by the network (Lempitsky et al., 2018). However, it has been observed that different classes have different statistics, and the relative size of buses is on average larger than that of cats. Consequently, we expect that different foveated retinotopies may emerge in different ecological niches.
In our case, it seems that the foveated retinotopic mapping allows for a more precise localization of the category of interest compared to off-the-shelf pre-trained networks using a Cartesian representation. It also gives us insight into the features on which our networks actually rely. Such information can be compared with physiological data (Crouzet, 2011), used to design better CNNs, and ultimately enable the development of physiological tests to further explore the features needed to classify a label of interest. In particular, by focusing on the point of fixation with the highest probability in the likelihood maps instead of the center of the bounding box defined in the “focus” dataset, we could consider refining the training of the network using our retinotopic mapping in a semi-supervised fashion.
@@
Building on these observations, simulating human-like saccadic eye movements offers a powerful approach to probe network mechanisms. A protocol that iteratively classifies foveated image patches could approximate natural viewing and reveal how performance varies across the visual field, as hypothesized in psychophysical experiments (Han et al., ). Comparing accuracy under different saccadic planning strategies—for example, prioritizing the most uncertain versus the most likely locations—would shed light on network attentional dynamics. This framework also enables modeling of biologically inspired saccade generation, allowing direct comparison with human visual search behavior.
Overall, implementing foveated classification with algorithmic saccades would provide a powerful method for validating existing attentional mechanisms in these networks, as well as inspiring new architectural innovations through embodied, task-driven visual attention modeling. In particular, this line of research would be especially suitable for dual-pathway models that effectively infer ego-motions (Bakhtiari et al., 2021; Mineault et al., 2021). Finally, the implementation of this robust categorization, coupled with refined localization of a label of interest and optimal saccade selection, could allow us to extend this study to more complex tasks. Visual search (i.e., the simultaneous localization and detection of a visual target) represents one such application where likelihood maps could provide the underlying pre-attentive mechanisms on which its effectiveness depends.
Materials and Methods
@@
In most mammals and amphibians, the arrangement of the (external) visual field is preserved in the early visual pathway, a feature called retinotopy. Retinotopic mapping results from the combined effect of the arrangement of photoreceptors in the retina and their output convergence via the optic nerve. This causes nearby regions of the visual field to activate adjacent neural structures as signals travel from the retina to the brain. These mappings differ from species to species, and our study concentrates on foveated vision (as in humans) which gives more resolution to the central field of view. In particular, we implement it by transforming the Cartesian pixel coordinates into log-polar coordinates (Javier Traver and Bernardino, 2010).
This simple parameterized transformation models accurately this biologically inspired retinotopic mapping. Considering arbitrary images (potentially with multiple channels such as RGB), each pixel’s position is defined by two real coordinates $(x,y)$ on a Cartesian reference frame. By convention, $x$ and $y$ are here considered belonging to the interval [1,1][-1,1], with $(0,0)$ being the center of the image. To implement the concentration of pixels near the center of the retina, we need to consider an irregular grid in the Cartesian referential that maps to a regular grid in the log-polar referential. In the log-polar referential, the location of each pixel has corresponding coordinates (${\rho},{\theta}$) as defined by previous studies (Hao et al., 2021) by:
@@
with $(x_{0},y_{0})$ defining the “center of fixation” (see Figure 1(Traver and Pla, 2003). Importantly, $\rho$ and $\theta$ are only defined for $(x,y)\neq(x_{0},y_{0})$. In most of our experiments, we consider $x_{0}=(0,0)$, allowing us to focus on the central part of the image. Each dot ($\rho$, $\theta$) in the log-polar coordinate system has thus a unique correspondence in the Cartesian coordinate system (and vice versa): For each ($\log\rho$, $\theta$) belonging to the grid, the corresponding pixel coordinate is $(x_{0}+\rho\cos\theta,y_{0}+\rho\sin\theta)$.
In practice, images have a finite resolution and to avoid biases in the evaluation between networks, the number of angles sampled ($N_{\theta}$) and the number of eccentricities sampled ($N_{\rho}$) are set to $224$, so that the size of the transformed image match the resolution $224\times 224$ of the input images. Note that native image resolution is generally higher than that used during network processing, with the average resolution in the Imagenet dataset being around $500\times 500$ pixels. This transformation is performed with the PyTorch library (Paszke et al., 2019) through the use of the grid_sample() function, which maps the pixels of an input image to the coordinates of any arbitrary grid, using a linear interpolation to estimate the value of the pixels. This function is used, for instance, in spatial transformer networks (Jaderberg et al., 2016).
@@
Convolutional neural networks (CNNs) have become essential tools in image classification, with several pre-trained models available for download. For example, the VGG family, including VGG16 and VGG19, introduced by Simonyan and Zisserman (2015), uses deep architectures with $16$ or $19$ layers, consisting of stacked convolutional layers followed by fully connected layers. ResNet (Residual Networks), introduced by He et al. (2015), addresses the vanishing gradient problem in deep networks by incorporating skip connections, allowing for the training of extremely deep networks (e.g. ResNet $50$, ResNet $101$ with respectively $50$ and $101$ layers). These models are widely used due to their robustness and scalability. The implementation of these deeper networks have demonstrated that deeper networks display enhanced resilience; however, this improvement is coupled with an overall increase in computational complexity (He et al., 2015). Therefore, based on these findings, we focus on the deep CNN ResNet (with $18$ to $101$ layers from the PyTorch library  [^1] We further introduced a circular padding in the convolutions, however we controlled this had little impact overall.) on the ImageNet-1K (Russakovsky et al., 2015) categorization challenge which consists in classifying natural images into $1000$ labels.
@@
Two datasets were used for our study: the first dataset is the one from the ImageNet (Russakovsky et al., 2015) challenge, which is the most widely benchmarked due to its huge collection of images and associated labels (the subset of ImageNet used in this study, i.e. ImageNet-1K with $1000$ labels and about $1000$ examples per label). In addition to the classification task, we consider here the localization task, which is prominent in computer vision. It consists in identifying both the label and the position of one or more objects of interest in an image. A distinction can be made between (i) the visual search task, where the label is given in advance and the task is simply to find the position of the object in the image, and (ii) the image labeling task, which consists in identifying both the objects and their position in the image, in the form of a labeled bounding box (see Figure 11). Despite its advantages, ImageNet has limitations for localization tasks. For example, the dataset lacks multi-labeling, with only one label per image. It is worth mentioning that ImageNet has some biases, the objects of interest are generally centered in the images, and the proportion of bounding boxes relative to the image size is relatively large, which may limit the impact of certain analyses.
@@
A widely accepted technique for evaluating the performance of Convolutional Neural Networks (CNNs) in localization tasks is the Class Activation Mapping (CAM) method. CAM works by analyzing the output of the CNN with respect to the target class, assigning weights to activations in each spatial feature map. This process generates a heat map that highlights significant areas of the image based on their contribution to the prediction. Building on the foundation of CAM, several derivative methods have emerged, including Grad-CAM (Selvaraju et al., 2020), Score-CAM (Wang et al., 2020), and Opti-CAM (Zhang et al., 2024).

### v5

@@ Results
We retrain pre-trained ResNet He2015wrn networks on different variants of the ImageNet-1K dataset Russakovsky2015 (see Methods “Datasets and Dataset Transformations”), either using a simple circular mask applied on the raw images (hereafter called the Cartesian dataset) or using log-polar transformed images (hereafter called the “retinotopic” dataset). Two variants of the training sets are also considered. In a first case (called the “regular” case), the mask or the log-polar transformed images are applied to the “regular” images. In a second case (called the “focus” case), the mask or the log-polar transformation is focused on the center of the bounding boxes surrounding the objects of interest and which are provided with the data set. The pre-trained networks are first re-trained on the “regular” version of the dataset, generating a first generation of networks, and then a second generation is trained by fine-tuning these networks on the “focus” dataset (see Figure 11). We then compare the classification accuracy of the original and the different re-trained networks. In this section we choose to express the accuracy as the percentage of correct categorization on the corresponding validation dataset.
@@
This difference between the two types of networks is even more manifest in Figure 2-B, where we compute the average accuracy over the test dataset for each single rotation angle. We first observe that the Cartesian networks show a decrease in accuracy with respect to the angle of rotation (with a symmetry with respect to horizontal flips), with a significant drop around $160^{\circ}$, “raw” and “focus” accuracies were degraded to $66.8\%$ and $68.1\%$, i.e. a drop of $16.1\%$ and $13.4\%$ ($64.3\%$ in “regular”, i.e. a drop of $15.8\%$). Strikingly, this effect is nearly absent in our retinotopic networks (see Figure 2-B), which show a flat (invariant) accuracy over the whole range of rotation angles, with a minimal degraded accuracy at $76.4\%$, i.e. a drop of $0.8\%$ ($71.6\%$ in “regular”, i.e. a drop of $2.9\%$). This marked difference can be interpreted as a consequence of the horizontal translation invariance found in classical CNNs. When applied to the retinotopic input space, this invariance transforms seamlessly into rotation invariance in visual space Araujo1997 (see Hao2021 for a proof). Note that for all networks, no maximum accuracy exceeds the value obtained without rotation (i.e. $0^{\circ}$ rotation angle) by more than $0.01\%$.
@@
To quantify the contribution of this sensitivity to translation, we consider here a new task, i.e. the visual search task, in which a visual object (of which the label is known in advance) needs to be localized over the entire image. Is is for instance known that such task allows for the quick retrieval of an image label Potter1975. We design a protocol for each network to allow us to compare different visual shifts, each one corresponding to a potential fixation point, and to generate a map of the expected (or actual) accuracy as a function of the fixation point (see Methods “Visual object localization : Protocol”). In this protocol, a set of $11\times 11$ fixation points is defined (with the coordinate $(5,5)$ being the center of the image), and at each coordinate of the grid, a likelihood value is computed for the label of interest (see figure 13). In practice, a likelihood is given at each location from a softmax calculation over the different labels, as in the classical ResNet classifier, providing a value between 0 and $1$ for the label of interest. This projects the network output onto a $11\times 11$ Bernoulli probability space corresponding to the likelihood of detecting the given label at each position, finally providing a “heat map” on our $11\times 11$ grid in a way that is compatible with other localization protocols. We tested our likelihood protocol on the “regular” validation dataset (see Fig 5 for some examples).
@@
ImageNet-1K provides rich semantic links that allow the construction of task-specific datasets. It has previously been demonstrated that the use of fine-tuning to re-train networks such as Vgg16 Simonyan2015 allows them to be applied to different tasks using the semantic network underlying ImageNet’s labels. Furthermore, it has been shown that the probability of a trained network performing a novel task (such as categorizing an animal) can be predicted using this semantic network, which links the outputs of a pre-trained network to a library of labels. This approach has been shown to be an effective method for learning to predict the presence of animals in images Jeremie2023. We can then exploit the semantic connection underlying the ImageNet labels by using these networks to perform a categorization task on the second dataset, the Animal 10k dataset, as referenced by Yu2021, with the goal of identifying the animals depicted in the images. This dataset contains $10,015$ images, each depicting an animal. All images were integrated in a validation dataset to test the networks. The output of each network is interpreted as “animal” if the label represented by the highest likelihood after softmax is indeed an animal. We use this prediction to calculate the accuracy of predicting an animal for all networks. Note that the accuracy is slightly biased due to the absence of distractors, it only represents the probability of finding an animal when there is actually an animal in the input. Figure 8 (right panel) compares the accuracy improvement for both networks. The general accuracy is high (around $95\%$ in all cases) due to our simpler binary (animal/non-animal) classification task. Saccade improvement is still significant, with all networks reaching a maximal accuracy of $100\%$ after a saccade to the maximum likelihood position for the label “animal”. The highest improvement is once again found in the retinotopic case (“raw” $4.5\%$, Cartesian $4.1\%$, retinotopic $6.3\%$).
@@ Discussion
The first and main result of this study is to demonstrate the excellent ability of standard deep CNNs to deal with foveated retinotopic inputs, even though this mapping enforces a radical transformation of the visual inputs. ResNet networks easily adapt to these inputs and the accuracy rates achieved with retinotopic inputs are equivalent to those of the original models. This is surprising given that the networks used in this re-training process were previously trained on Cartesian images. Images with a log-polar transformation show a high degree of distortion, particularly a high compression of visual information around the fixation point and a degradation of textures in the periphery (see Figure 1). One possible hypothesis is that the degradation of texture during the frame of reference change may cause the network to rely on shape rather than texture da_costa_convolutional_2024. A further study could entail a comparison of these visual processing parameters with those of human vision, thereby providing insight into their evolutionary trajectory.
@@
In addition, the log-polar transformation has the advantage of better invariance to zooms and rotations. First, this study shows that for the original ResNet network on “regular” images, the average accuracy with a zoom or a rotation dropped sharply compared to baseline accuracy, confirming that these simple geometrical transformations misled the networks. For rotations, the decline was steepest around $160^{\circ}$, demonstrating limited rotational invariance compared to humans Rousselet2003; Guyonneau2006. The robustness differs slightly when studying the impact of zooms. This may be attributed to the transformation under examination, which already provides a zooming-in effect of the fovea that accentuates the information in the area surrounding the point of fixation. To test this hypothesis, we could train a retinotopic version without applying the logarithmic function to the eccentricity axis. Finally, this invariance to rotations and zooms comes at the cost of reduced invariance to translations. For images not centered on the region of interest, one would need to shift the fixation point to the region of interest, similar to eye saccades in biological vision. The effects of this trade-off can be observed in the comparison between the “regular” version of the networks and the “focus” version.
@@
In addition, the log–polar transformation confers improved invariance to zoom and rotation. In the original ResNet trained on “regular” Cartesian images, accuracy declined sharply under zooms or rotations, confirming that even simple geometric transformations can mislead the network. The steepest drop occurred around $160^{\circ}$, highlighting limited rotational invariance compared to humans Rousselet2003; Guyonneau2006. Robustness to zoom was less affected, likely because the transformation itself magnifies foveal input relative to the periphery. This robustness, however, comes at the cost of reduced tolerance to translations: images misaligned with the fixation point require shifting the sampling center, analogous to saccades in biological vision. This trade-off is evident when comparing the “regular” and “focus” network variants. Similar effects have been reported previously:  pun_log-polar_2003 showed that log–polar wavelet signatures enhance joint scale–rotation invariance, while  amorim_analysing_2018 demonstrated improved rotational robustness in natural image categorization and highlighted the positional sensitivity introduced by inhomogeneous log–polar sampling. Finally, previous research has demonstrated that improved invariance to scale variations is often associated with greater robustness to input compression Remmelzwaal2019. This dual robustness — the ability to maintain performance across scaled inputs and compressed representations — facilitates the application of categorisation models to high-resolution images without compromising accuracy. These properties are particularly valuable in computational neuroscience, where visual processing systems must generalise across diverse input conditions.
@@
One hypothesis is that the foveated retinotopic mapping implemented in the log-polar transform provides an efficient prior for visual object geometry. Indeed, a main source of variability in the view of an object comes from displacements of the observer relative to it, for example as the object scales in the visual field as one approaches it, or as the object rotates with a rotation of the head. The log-polar mapping allows for a more invariant representation which explicitly implements this prior for these displacements. Additionally, imposing translation invariance of the representation establishes a prior on the possible representations allowed by the network Lempitsky2018. However, it has been observed that different classes have different statistics, and the relative size of buses is on average larger than that of cats. Consequently, we expect that different foveated retinotopies may emerge in different ecological niches.
In our case, it seems that the foveated retinotopic mapping allows for a more precise localization of the category of interest compared to off-the-shelf pre-trained networks using a Cartesian representation. It also gives us insight into the features on which our networks actually rely. Such information can be compared with physiological data Crouzet2011, used to design better CNNs, and ultimately enable the development of physiological tests to further explore the features needed to classify a label of interest. In particular, by focusing on the point of fixation with the highest probability in the likelihood maps instead of the center of the bounding box defined in the “focus” dataset, we could consider refining the training of the network using our retinotopic mapping in a semi-supervised fashion.
@@
Building on these observations, simulating human-like saccadic eye movements offers a powerful approach to probe network mechanisms. A protocol that iteratively classifies foveated image patches could approximate natural viewing and reveal how performance varies across the visual field, as hypothesized in psychophysical experiments han_scale_2020. Comparing accuracy under different saccadic planning strategies—for example, prioritizing the most uncertain versus the most likely locations—would shed light on network attentional dynamics. This framework also enables modeling of biologically inspired saccade generation, allowing direct comparison with human visual search behavior.
Overall, implementing foveated classification with algorithmic saccades would provide a powerful method for validating existing attentional mechanisms in these networks, as well as inspiring new architectural innovations through embodied, task-driven visual attention modeling. In particular, this line of research would be especially suitable for dual-pathway models that effectively infer ego-motions Bakhtiari2021; Mineault2021. Finally, the implementation of this robust categorization, coupled with refined localization of a label of interest and optimal saccade selection, could allow us to extend this study to more complex tasks. Visual search (i.e., the simultaneous localization and detection of a visual target) represents one such application where likelihood maps could provide the underlying pre-attentive mechanisms on which its effectiveness depends.
Methods
@@
In most mammals and amphibians, the arrangement of the (external) visual field is preserved in the early visual pathway, a feature called retinotopy. Retinotopic mapping results from the combined effect of the arrangement of photoreceptors in the retina and their output convergence via the optic nerve. This causes nearby regions of the visual field to activate adjacent neural structures as signals travel from the retina to the brain. These mappings differ from species to species, and our study concentrates on foveated vision (as in humans) which gives more resolution to the central field of view. In particular, we implement it by transforming the Cartesian pixel coordinates into log-polar coordinates JavierTraver2010.
This simple parameterized transformation models accurately this biologically inspired retinotopic mapping. Considering arbitrary images (potentially with multiple channels such as RGB), each pixel’s position is defined by two real coordinates $(x,y)$ on a Cartesian reference frame. By convention, $x$ and $y$ are here considered belonging to the interval [1,1][-1,1], with $(0,0)$ being the center of the image. To implement the concentration of pixels near the center of the retina, we need to consider an irregular grid in the Cartesian referential that maps to a regular grid in the log-polar referential. In the log-polar referential, the location of each pixel has corresponding coordinates (${\rho},{\theta}$) as defined by previous studies Hao2021 by:
@@
with $(x_{0},y_{0})$ defining the “center of fixation” (see Figure 1Traver2003. Importantly, $\rho$ and $\theta$ are only defined for $(x,y)\neq(x_{0},y_{0})$. In most of our experiments, we consider $x_{0}=(0,0)$, allowing us to focus on the central part of the image. Each dot ($\rho$, $\theta$) in the log-polar coordinate system has thus a unique correspondence in the Cartesian coordinate system (and vice versa): For each ($\log\rho$, $\theta$) belonging to the grid, the corresponding pixel coordinate is $(x_{0}+\rho\cos\theta,y_{0}+\rho\sin\theta)$.
In practice, images have a finite resolution and to avoid biases in the evaluation between networks, the number of angles sampled ($N_{\theta}$) and the number of eccentricities sampled ($N_{\rho}$) are set to $224$, so that the size of the transformed image match the resolution $224\times 224$ of the input images. Note that native image resolution is generally higher than that used during network processing, with the average resolution in the Imagenet dataset being around $500\times 500$ pixels. This transformation is performed with the PyTorch library Paszke2019 through the use of the grid_sample() function, which maps the pixels of an input image to the coordinates of any arbitrary grid, using a linear interpolation to estimate the value of the pixels. This function is used, for instance, in spatial transformer networks Jaderberg2016.
@@
Convolutional neural networks (CNNs) have become essential tools in image classification, with several pre-trained models available for download. For example, the VGG family, including VGG16 and VGG19, introduced by Simonyan2015, uses deep architectures with $16$ or $19$ layers, consisting of stacked convolutional layers followed by fully connected layers. ResNet (Residual Networks), introduced by He2015wrn, addresses the vanishing gradient problem in deep networks by incorporating skip connections, allowing for the training of extremely deep networks (e.g. ResNet $50$, ResNet $101$ with respectively $50$ and $101$ layers). These models are widely used due to their robustness and scalability. The implementation of these deeper networks have demonstrated that deeper networks display enhanced resilience; however, this improvement is coupled with an overall increase in computational complexity He2015wrn. Therefore, based on these findings, we focus on the deep CNN ResNet (with $18$ to $101$ layers from the PyTorch library on the ImageNet-1K Russakovsky2015 categorization challenge which consists in classifying natural images into $1000$ labels. We further introduced a circular padding in the convolutions, however we controlled this had little impact overall.
@@
Two datasets were used for our study: the first dataset is the one from the ImageNet Russakovsky2015 challenge, which is the most widely benchmarked due to its huge collection of images and associated labels (the subset of ImageNet used in this study, i.e. ImageNet-1K with $1000$ labels and about $1000$ examples per label). In addition to the classification task, we consider here the localization task, which is prominent in computer vision. It consists in identifying both the label and the position of one or more objects of interest in an image. A distinction can be made between (i) the visual search task, where the label is given in advance and the task is simply to find the position of the object in the image, and (ii) the image labeling task, which consists in identifying both the objects and their position in the image, in the form of a labeled bounding box (see Figure 11). Despite its advantages, ImageNet has limitations for localization tasks. For example, the dataset lacks multi-labeling, with only one label per image. It is worth mentioning that ImageNet has some biases, the objects of interest are generally centered in the images, and the proportion of bounding boxes relative to the image size is relatively large, which may limit the impact of certain analyses.
@@
A widely accepted technique for evaluating the performance of Convolutional Neural Networks (CNNs) in localization tasks is the Class Activation Mapping (CAM) method. CAM works by analyzing the output of the CNN with respect to the target class, assigning weights to activations in each spatial feature map. This process generates a heat map that highlights significant areas of the image based on their contribution to the prediction. Building on the foundation of CAM, several derivative methods have emerged, including Grad-CAM Selvaraju2020, Score-CAM Wang2020, and Opti-CAM Zhang2024.
@@
Declarations
Data Availability
All results presented here are based on data that are openly available, as described in the Methods section.
Code Availability
The code to reproduce all results and figures is available at https://github.com/bicv/Retinotopy.